From GlueXWiki
Jump to: navigation, search

July 12, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • Fourth package tests
    • Spare wire frames
    • Third package preparation
  2. Engineering update (Bill)
  3. Electronics update: PR for the spare PCBs, grounding, other (Chris, Nick)
  4. Gas system update (Beni)
  5. Test set-up in 126 (Beni)
  6. Other


Participants: Bill, Dave, Eugene, Mark, Nick, Simon, Vlad, and Lubomir


- Results of the tests of the last three cells of the fourth package: top cell #1 (wire frame #18) working fine, cell #2 (wire frame #28) and cell #3 (wire frame #27) had a problem, high dark currents, about 400-500nA for frame #28 and 1100-1300nA for #27. Both cells held HV for a long time (3-4 days) and all channels are OK. Still, we were not comfortable with that and disassembled the three cells on Monday. It turned out that the HV caps on all wire frames were indeed replaced, just Anatoly didn't have record for #28. We can't find what was different in the production of the last two wire frames. On Wednesday we installed back the third cell but with the working wire frame #18, that was before on the top.

- From Monday till Wednesday we were testing #27 separately between the two Lexan sheets: the high current didn't change even after cleaning, or removing HV cap on the HV PCBs. There's no current through the field HV and looks like the current goes on the HV PCBs to ground. Most likely humidity plays a role here. We used a lamp for several hours to heat/dry the PCB; the current goes up by a factor of two almost immediately after turning it on. On Wednesday we gave up with #27 and started testing wire frame #28, the current there is much lower, ~400nA. The plan is to try to clean it and if it's stable we will use it in the second cell. As for #27, the plan for now is to leave it as a spare and use one of the spare frames (#6) in the fourth package.

- Spare wire frame #6: Anatoly was able to remove the HV caps and clean the Humiseal (without using aceton); there's die in the Humiseal which makes it very easy to see it with UV light. Next, we will test the wire frame between the Lexan sheets, first, without HV caps. Since we are not sure if we can fix #27, Bill will ask "Vision Machine" for quotation of four g10 wire frames. If we need them, we have to submit PR before the end of July.

- Third package: 10 ground clips (first article) finally arrived and we tested them on the package. They have a bite of ~12.5mm and we asked for 11mm, that's why we have to bend them a little, but other than that they worked fine. Casey put already them on one cell and is looking for a way to bend them in a standard way, using a press. This is not a big deal and we decided to go ahead with the production of the rest of the clips, instead of waiting another month. Tina and Dave put the labels on all the pre-amps for the third package.

- We decided to try a new procedure for the installation of the cooling tubes, as proposed by Casey. Initially we don't put glue on the back brackets, just screw them loosely to the pre-amp brackets with the tube in between. After adjusting the tube position we tight the screws and then put glue on the back brackets from the sides. Mark suggested to tight the two ends of the tube together, before, on the first package, we left them loose.

Engineering update

- Bill went through the 3D models of the cooling system attached above, explaining part by part both the water loop and the Fluorinert loop. He tries to find a way to detect early possible leakage. It was suggested to install a camera at the cooling system cart.

Electronics update

- Nick tested the LV distribution board, had to put 6.4V on the power supply to get at least 3V on the cards. Eugene had two remarks: one is that we double the power, will have the same heat on the LV cables as on the pre-amps. Bill: since only a quoter of the cables is inside the magnet, the cooling system will still be able to take the heat out of the detectors. Second Eugene's remark: we have four pre-amps on one cable and what if three of them fail, the current will be a factor of 4 lower, therefore the voltage on the last pre-amp will go up ~5.5V, is that OK for the voltage regulator?

- PR for the extra PCBs. Eugene: Fernando is aware of the problem with the sole source and that it has to be submitted by the end of the month.

- Grounding: since we use copper foil on coper/gold we discussed if the copper oxidation will affect the grounding. Bill: according to Fernando copper oxide is conductive. Just in case there are problems we can solder the foil to the cathode/PCB surfaces.

Gas system update

- There was a meeting on Wednesday, deciding to make budget estimates for the different options and see how it will fit in the budget. Two aspects directly related to FDC were discussed here:

- The possibility to have a third gas in the mixture. We may need it in case our resolution is not good enough and we need to increase the gain (previous studies suggested the resolution improves with the gain). Then we may have some after-pulsing that can be suppressed with ternary gas mixture. Apart from the price Eugene said, first we want to make sure the use of organic gas doesn't result in aging. Lubomir: aging is proportional to the rates and as an example drift chambers with almost the same wires were used in the BigBite spectrometer in Hall A for many years at much higher rates. However, they used ethane, not isobutane. Bill: what if we pre-mix our gas somewhere else, what is the precision for the isobutane percentage? Lubomir will prepare a presentation on the gas mixtures for the next meeting.

- Everybody agreed that we need to monitor the pressure of each individual cell using the gauge tubes. Bill: in case of some pressure difference between the cells we can use the same tubes to regulate the pressure in some extent.

Test set-up at 126

- According to Beni the DAQ is almost ready, some problems with reading some of the modules, but the main problems have been solved.